OpenStudy (anonymous):

Simplify: (7x + 1) (7x - 1)

OpenStudy (anonymous):

Use disribution twice. Start with (7x + 1) (7x - 1) = 7x(7x-1) + 1(7x-1)

OpenStudy (abb0t):

\[(7x \times 7x)+(7x \times -1)+(7x \times 1) + (1 \times -1)\]